Please answer using grade 11 Conservation of Energy and Momentum please!

Thank-you!

You and a colleague are on a spacewalk, repairing your spacecraft that has stalled in deep space. Your 90.0 kn colleague, initially at rest, asks you to throw her a hammer, which has a mass of 3.0 kg. You throw it to her with a velocity of 4.5 m/s [forward].

a) What is her velocity after catching the hammer?

b) What impulse does the hammer exert on her?

c) What percentage of kinetic energy is lost in the collision?
1
Expert's answer
2020-05-12T09:32:28-0400

(a)initial momentum = final momentum

So


"3\times4.5= (3+90)\ v\\or\\v=0.1452\ m/s"

(b) impulse = change in momentum of hammer=

"m(v-4.5)=3(0.1452-4.5)=13. 0644\ kg-m/s"

(C) initial kinetic energy =

"\frac{1}{2}m\times4.5^2=0.5\times 3\times 20.25=30.375\ J"

Final kinetic energy


"\frac{1}{2}(90+3)(0.1452)=6.7518\ J"

Percentage loss in energy = "\frac{30.375-6.7518}{30.375}\times100= 77
.77\ percent"